dP/dT+2tP=P+4t-2


1
Expert's answer
2021-01-26T18:52:33-0500

"dp\/dt+2tp=p+4t-2"

"dp\/dt=p+4t-2-2tp=(2-p)(-1+2t)"

"dp\/dt=(2-p)(-1+2t)"

"\int dp/(2-p)=\int(-1+2t)dt"

"-ln (2-p)=-t+t^2+C_3"

"2-p=exp(t-t^2+C_2)=C_1exp(t-t^2), C_2=-C_3, C_1=exp(C_2)"

"p(t)=Cexp(t-t^2)+2, C=-C_1"